Ka for ethanoic acid is 1.7 × 10–5 mol dm–3 at 25 °C.

What is the pH of a 0.125 mol dm–3 solution of ethanoic acid at this temperature?


1
Expert's answer
2022-05-06T08:37:04-0400

Solution:

ethanoic acid ≡ acetic acid ≡ HC2H3O2

Ka for ethanoic acid is 1.7×10–5 mol dm–3


The balanced equation for the ionization of HC2H3O2 is

HC2H3O2(aq) + H2O(l) ⇌ C2H3O2(aq) + H3O+(aq)

Summarize the initial conditions, the changes, and the equilibrium conditions in the following ICE table:



Substitute the equilibrium concentrations into the expression for the acid ionization constant Ka:

Ka = [C2H3O2] × [H3O+] / [HC2H3O2] = (x) × (x) / (0.125 − x)


Make the approximation that 'x' is small (0.125 >> x), substitute the value of the acid ionization constant into the Ka expression and solve for 'x':

1.7×10–5 = (x2) / (0.125)

x2 = 2.125×10–6

x = 1.458×10–3


The ratio of 'x' to the number it was subtracted from and added to in the approximation is less than 0.05; therefore, the 'x' is small approximation is valid, and [H3O+] = x = 1.458×10–3 M.


Finally, calculate the pH as the negative of the base-10 logarithm of [H3O+]:

pH = −log[H3O+] = −log(1.458×10–3) = 2.836 = 2.84

pH = 2.84


Answer: The pH of ethanoic acid solution is 2.84